OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. sine of x divided by one minus cosine of x + sine of x divided by one minus cosine of x = 2 csc x

OpenStudy (anonymous):

\[\sin x/ 1 - \cos x + \sin x / 1 + \cos x = 2 \csc x\]

OpenStudy (anonymous):

please helppp

OpenStudy (zale101):

\[\frac{ sinx }{ 1-cosx }+\frac{ sinx }{ 1+cosx }=2cosx\] Does your equation looks similar to this?

OpenStudy (zale101):

2cscx *